# RobustMerge: Parameter-Efficient Model Merging for MLLMs with Direction Robustness

### v2

@@ Appendix C More Implementation Details
We build multimodal codebase for multimodal tasks and vision tasks upon LLaVA  [^2] https://github.com/haotian-liu/LLaVA and CLIP  [^3] https://github.com/openai/CLIP, respectively. For training, we follow the standard training procedure described in LLaVA, i.e., training each task individually and obtaining parameter-efficient modules. LoRA is added to linear layers in foundational blocks and all models are trained for 1 epoch for merging.
For inference, the tasks are evaluated by accuracy. Vision task is conducted by constructing textual prompts for each category and calculating the similarity. Each evaluated sample is classified into the category with the largest similarity. Hyper-parameter $\lambda$ is set to 2 by default. All merging experiments are carried out on a single NVIDIA A6000 with the temperature set to 0.
@@ Appendix D Details of Comparison Methods
The crucial operation of merging method involves designing a merging algorithm, i.e., $\Phi(\cdot)$ defined in the paper. We primarily compare with DARE, Ties-merging, and Task Arithmetic in the main results. For comparison to these traditional approaches, we employ traditional approaches to LoRA components. Specifically, we fine-tune foundation model using LoRA, merge LoRA components using traditional approaches and finally evaluate performance with merged LoRA attached to base models. Their ways of merging are briefly introduced in the following paragraph.
Task Arithmetic views all parameters as vectors. It obtains task vectors by subtracting fine-tuned models and pre-trained models and performs standard arithmetic operations like addition and subtraction on them. It sums the parameters from checkpoints of different tasks and constructs a strong baseline for multi-task learning.
Ties-merging reduces the conflicts between parameters of different tasks by a trim, elect sign and merging paradigm. Concretely, it first keeps parameters with the highest magnitudes, and then determines the aggregated sign based on the summation of remaining parameters. Finally, it merges the parameters with the same sign as the aggregated sign to mitigate disagreements.
@@ E.2 Datasets of Vision Tasks
All vision tasks are traditional datasets containing common objects across wide domains like cars, texture, traffic signs and so on for image classification. Categories for them vary from 10 to 397. We fine-tune VLMs with LoRA on each task and merge them employing different types of merging methods. Only the visual encoder is trainable and the text encoder remains frozen for label embedding extraction.
@@ Appendix F More Experimental Results
RobustMerge generalizes to the number of tasks. We gradually increase the number of tasks to substantiate the robustness of our method. As is illustrated in Fig. 13, in seen tasks, the performance undergoes a slight drop as merging more models interferes specific task. In unseen tasks, the performance first improves and then declines modestly. It could be attributed that in the first stage, seen tasks transfer knowledge and enhance unseen tasks of similar distribution; In the second stage, interference dominates merging rather than knowledge transformation. Under both circumstances, our method consistently outperforms existing approaches by a substantial margin as the task number varies, indicating the superiority and stability of our method.
@@
RobustMerge extends well to different PEFT methods. We primarily test our method on LoRA since it is the most commonly used and comparable PEFT techniques. To demonstrate the extensibility of the proposed method, we additionally conduct experiments DoRA [32], which is LoRA-based efficient technique with an advanced algorithm to improve the performance of PEFT learning. Results shown in Tab. 5 reveal that our method achieves consistent and substantial improvements against existing merging methods in different PEFT methods. It strongly certifies that the problem of direction robustness is widespread in merging different kinds of PEFT modules, where attention is primarily paid to improving the performance of a single task. By contrast, our method handles the issue to some extent, thereby promoting the multi-task performance when the PEFT technique varies.

### v5

@@ Appendix C More Implementation Details
We build a multimodal codebase for multimodal tasks and vision tasks upon LLaVA  [^2] https://github.com/haotian-liu/LLaVA and CLIP  [^3] https://github.com/openai/CLIP, respectively. For training, we follow the standard training procedure described in LLaVA, i.e., training each task individually and obtaining parameter-efficient modules. LoRA is added to linear layers in foundational blocks, and all models are trained for 1 epoch for merging.
For inference, the tasks are evaluated by accuracy. The vision task is conducted by constructing textual prompts for each category and calculating the similarity. Each evaluated sample is classified into the category with the largest similarity. Hyper-parameter $\lambda$ is set to 2 by default. All merging experiments are carried out on a single NVIDIA A6000 with the temperature set to 0.
@@ Appendix D Details of Comparison Methods
The crucial operation of merging method involves designing a merging algorithm, i.e., $\Phi(\cdot)$ defined in the paper. We primarily compare with DARE, Ties-merging, and Task Arithmetic in the main results. For comparison to these traditional approaches, we employ traditional approaches to LoRA components. Specifically, we fine-tune the foundation model using LoRA, merge LoRA components using traditional approaches, and finally evaluate performance with merged LoRA attached to base models. Their ways of merging are briefly introduced in the following paragraph.
Task Arithmetic views all parameters as vectors. It obtains task vectors by subtracting pre-trained models from fine-tuned models and performs standard arithmetic operations like addition and subtraction on them. It sums the parameters from checkpoints of different tasks and constructs a strong baseline for multi-task learning.
Ties-merging reduces the conflicts between parameters of different tasks by a trim, elect sign, and merging paradigm. Concretely, it first keeps parameters with the highest magnitudes, and then determines the aggregated sign based on the summation of remaining parameters. Finally, it merges the parameters with the same sign as the aggregated sign to mitigate disagreements.
@@ E.2 Datasets of Vision Tasks
All the vision tasks are traditional datasets containing common objects across wide domains like cars, texture, traffic signs, and so on for image classification. Categories for them vary from 10 to 397. We fine-tune VLMs with LoRA on each task and merge them employing different types of merging methods. Only the visual encoder is trainable, and the text encoder remains frozen for label embedding extraction.
@@ Appendix F More Experimental Results
RobustMerge generalizes to the number of tasks. We gradually increase the number of tasks to substantiate the robustness of our method. As is illustrated in Fig. 13, in seen tasks, the performance undergoes a slight drop as merging more models interferes with specific tasks. In unseen tasks, the performance first improves and then declines modestly. It could be attributed to the fact that in the first stage, seen tasks transfer knowledge and enhance unseen tasks of similar distribution; in the second stage, interference dominates merging rather than knowledge transformation. Under both circumstances, our method consistently outperforms existing approaches by a substantial margin as the task number varies, indicating the superiority and stability of our method.
@@
RobustMerge extends well to different PEFT methods. We primarily test our method on LoRA since it is the most commonly used and comparable PEFT technique. To demonstrate the extensibility of the proposed method, we additionally conduct experiments on DoRA [34], which is a LoRA-based efficient technique with an advanced algorithm to improve the performance of PEFT learning. Results shown in Tab. 5 reveal that our method achieves consistent and substantial improvements against existing merging methods in different PEFT methods. It strongly certifies that the problem of direction robustness is widespread in merging different kinds of PEFT modules, where attention is primarily paid to improving the performance of a single task. By contrast, our method handles the issue to some extent, thereby promoting the multi-task performance when the PEFT technique varies.